# Influence of the Galactic Halo on the UHECR Multipoles

### v1

@@ 1 Introduction
The arrival directions of ultra-high energy cosmic rays (UHECRs) at Earth is not entirely isotropic. In 2017, the Pierre Auger Collaboration (PAO) reported the detection of dipolar anisotropy, for UHECRs with energies $\geq 8$ EeV (Aab et al., 2017). This result was updated in Refs. Aab et al. (2018b) and Abdul Halim et al. (2024) and is now at a significance of $6.8\sigma$ with a dipole amplitude of $\delta=\rm{0.074}^{+0.010}_{-0.008}$ being obtained. The direction of this dipole lies ${\sim}115\degree$ away from the Galactic centre, suggesting an extragalactic UHECR origin. It is worth noting that this cosmic ray dipole does not align with the CMB dipole or any potential known source.
The Telescope Array collaboration (TA) also searched for a large-scale dipole anisotropy in the northern hemisphere sky (Abbasi et al., 2020). They reported a dipole amplitude of $3.3\pm 1.9\%$ which, at their current level of statistics, is compatible with the dipole reported by PAO, but also with an isotropic distribution. When the datasets of PAO and TA are combined, a dipole with a significance of ${\sim}4.2\sigma$ is found (Abdul Halim et al., 2023), compatible with the dipole reported by PAO alone. Besides the dipole anisotropies, several indications have been found of intermediate-scale anisotropies at higher energies as well (see e.g. Refs. Abbasi et al. (2018); Abreu et al. (2022)). However, our focus in this paper is predominantly on the large-scale anisotropy.
@@
The observed UHECR dipole strength at Earth for cosmic rays with energies above 8 EeV is at the 7% level. However, this dipole strength is not necessarily the same as the arriving extragalactic dipole amplitude, out at large distances where the influence of the Milky Way’s magnetic fields has yet to take effect. The Galactic magnetic fields can affect an external extragalactic dipole ($\delta_{\rm{EG}}$) in two main ways: the large-scale structured component changes the extragalactic direction of $\delta_{\rm{EG}}$ and the small-scale turbulent magnetic fields (with RMS value of $B_{\rm{tur}}$) suppress the strength of $\delta_{\rm{EG}}$ (Zirakashvili, 2005). Furthermore, these effects suggest that a beam (current) of particles from a source, upon propagating through the Galactic magnetic field structure, is also expected to be changed, potentially with the arrival directions being spread out by the turbulent field component, resulting in a distribution resembling a dipole.
The ability of the Milky Way’s turbulent magnetic field to alter the cosmic ray distribution locally observed is dependent on the spatial extent and turbulence level of this field component. Over the last few years, a growing body of evidence has accumulated indicating that a considerable level of thermal gas pressure exists out in the halo of the Milky-Way and other local galaxies of a similar mass, out to at least a distance of the virial radius (see Martynenko (2022); Bregman et al. (2022); Zhang et al. (2024)). Assuming that the plasma beta of this halo gas is close to unity, a considerable magnetic field strength is also expected out in this region Faerman et al. (2020). Interestingly, such an inference is consistent with recent radio observations of the circumgalactic medium region (at a distance $\sim 100$ kpc) of cosmologically nearby Galaxies. From such studies, the line-of-sight magnetic field in these regions have been found to have strengths of a few 10s of % of $\mu$G, using Faraday rotation measurements (Heesen et al., 2023).
In this work we build on these recent results by considering the effects of an extended turbulent Galactic magnetic field halo on the penetration of cosmic rays to the local observer. In particular, we investigate the effect of the GMF on the arrival of an extragalactic UHECR large-scale dipole distribution, or an extragalactic UHECR small-scale localized distribution. Our focus in this paper is on small-scale turbulent GMFs, as those will have the strongest effect on the expected amplitude of the dipole and quadrupole.
@@ 2 Methodology - Simulation setup
UHECRs gyrating with a Larmor radius larger than the coherence length of the magnetic field, ie. $r_{L}\gg L_{\rm coh}$, in isotropic turbulent magnetic fields, undergo only small angular deflections of size $\Delta_{\Theta}=L_{\rm{coh}}/r_{L}$ in each coherent magnetic field patch, where $\Theta$ is the angle between the particle’s initial and final momentum direction. For a particle propagating a path length $L$, the number of coherence cells traversed, $N$, is $N=L/L_{\rm{coh}}$. The standard deviation in the deflection angles from this number of small-scale deflections is,  $\sigma_{\Theta}=\sqrt{N}\Delta_{\Theta}$. After a sufficient number of scatterings, the spread in angles becomes large, $\sigma_{\Theta}\sim\pi$ radians. The large angle scattering length can therefore be estimated to be, $L_{\rm{scat}}\approx(\pi/\Delta_{\Theta})^{2}L_{\rm coh}=\pi^{2}\frac{r_{L}^{2}}{L_{\rm{coh}}}$. The diffusion coefficient $D\sim L_{\rm{scat}}~{}\rm{c}/3$, where ’$\rm{c}$’ is the speed of light in a vacuum. Throughout the text will use the terms ”scattering length” and ”diffusion coefficient” interchangeably.
We use CRPropa 3 (Alves Batista et al., 2016, 2022) to generate isotropic Kolmogorov 5/3 turbulent magnetic fields with $L_{\rm min}$ = 200 pc and $L_{\rm max}$ = 400 pc as the minimum and maximum wavelengths respectively. We stick to this restricted dynamic range of $L_{\rm min}$ and $L_{\rm max}$ for computational reasons (see Shaw et al. (2022) for more details about the field). With the coherence length of the field fixed at 150 pc, the turbulent field effectively has only one free parameter, the magnitude of the turbulent field strength, $B_{\rm tur}$. While large, this coherence length value, $L_{\rm coh}$, falls within ranges previously studied in the literature (Ohno & Shibata, 1993; Chepurnov et al., 2010; Iacobelli et al., 2013; Beck et al., 2016; Giacinti et al., 2018). However, it should be borne in mind that these other studies predominantly focussed on more central Galactic distances than those considered here, and that the actual coherence length could, therefore, be considerably larger than that adopted here. The turbulent fields in our setup extend out to the simulation absorption boundary (termination sphere) with radius $R_{\rm T}=60$ kpc (see Fig. 1).
We note that the length scale values chosen for our setup, specifically $R_{\rm T}$ and $R_{\rm I}$, are chosen for computational rather than physical reasons. The purpose or our work, instead, is a proof of principle of the shielding effect produced by a magnetised halo, rather than an accurate description of the halo.
@@ 2.1 Estimation of the scattering length
To determine the scattering length ($L_{\rm{scat}}$) of the UHECRs from simulations, we inject these particles in a single direction into the turbulent magnetic field setup described in Sec. 2. For each value of $B_{\rm{tur}}$, we ran the simulation until the particles had completely isotropised such that $\langle|{\Theta}|\rangle$, had plateaued at $\approx\pi/2$. The trajectory length at which $\langle|{\Theta}|\rangle$ plateaus thereby signifies that the particles have lost memory of their injected angular distribution, and is referred to as the particle scattering length $L_{\rm{scat}}$. In order to determine $L_{\rm{scat}}$ for different values of $B_{\rm{tur}}$, we vary $B_{\rm{tur}}$ across 15 logarithmically spaced values from 2 to 35 $\mu$G, whilst maintaining the same coherence length, $L_{\rm{coh}}=150$ pc. Test simulations were performed for several coherence length values, but for consistency with Shaw et al. (2022), the 150 pc value was chosen.
Figure 1: Schematic representation of the simulation setup for the dipole and beam setup shown on the left and right-hand side respectively. In both figures, $R_{\rm T}$, $R_{\rm O}$ and $R_{\rm I}$ are the radii of the termination, observer and injection spheres respectively. The distribution of the red arrows in the dipole setup indicates the injected external dipole and the single arrow in the beam setup denotes injected UHECRs with the arrow directions pointing inwards in both setups.
Comparing these numerical results with our earlier analytic expectations, for small values of $B_{\rm{tur}}$, such that $r_{L}\gg L_{\rm coh}$, the relation $L_{\rm{scat}}\approx\pi^{2}\frac{r_{L}^{2}}{L_{\rm{coh}}}$ appears to hold (see Fig. 2). This changes however, as the magnetic field strength further increases and the scattering length of the particle approaches the turbulence coherence length scale i.e. the particle is more subjected to large-angle scattering. For such cases, we see a deviation from the dependence $L_{\rm{scat}}\approx\pi^{2}\frac{r_{L}^{2}}{L_{\rm{coh}}}$.
Figure 2: The scattering length plotted as a function of increasing turbulent magnetic field strength, for 8.5 EeV protons and a turbulence coherence length of $150$ pc. The dashed line shows the expected results applicable when $r_{L}\gg L_{\rm coh}$.
@@ 2.2 Calculation of UHECR dipole amplitudes through turbulent magnetic fields
In order to create synthetic skymaps for an observer we inject between $10^{9}-10^{10}$ UHECRs into the setup depicted in Fig. 1. This setup ensures that the observed skymaps contain good statistics ($4\times 10^{5}$). In reality, UHECRs enter the Galaxy from all directions, and what is observed by PAO (Aab et al., 2017; Abreu et al., 2022) is a fraction of the UHECRs that hit the observer (Earth).
Figure 3: Top-left: panel shows the injected extragalactic dipole skymap. Top-right: shows the single-beam injected skymap. Bottom-left: panel depicts the combined three-beam injected skymap. Bottom-right: panel shows the comparative power spectrum analysis of the three injected skymaps. Note that all maps have a $10^{\circ}$ top-hat smoothing applied. This is similar to what is used by PAO (Abdul Halim et al., 2024).
Forward tracking of UHECRs was performed using CRPropa 3, where 8.5 EeV protons were propagated through the Galaxy’s turbulent magnetic field component after being injected inward from an injection sphere. The observer sphere radius $R_{\rm O}$ was set to 150 pc, comparable to the size of $L_{\rm{coh}}$. The simulation ends when the particle reaches either the observer sphere or the termination sphere (see Fig. 1). Although for an ideal simulation, the size of the observer should be as small as possible to mimic the Earth as an observer sphere, due to computational limitations we choose an observer sphere size of 150 pc. The injection sphere was kept at a radius, $R_{\rm I}=$ 30 kpc and the escape boundary was twice as large as the injection sphere at $R_{\rm T}=$ 60 kpc.
@@
    An extragalactic dipole case where we impose an external dipole at $\sqrt{3}$ level at the injection sphere (see top-left skymap Fig. 3). The value $\sqrt{3}$ comes from the prefactor in $\delta_{0}=\sqrt{3}\frac{C_{1}}{C_{0}}$.
@@
    A three beam case where we inject three $10^{\circ}$ beams of UHECRs at the positions of Cen A, NGC 253 ($l=97.4^{\circ}$, $b=-87.9^{\circ}$) and M82 ($l=141.4^{\circ}$, $b=+40.6^{\circ}$) (see bottom-left skymap Fig. 3). In the three-beam configuration, two of the beam directions, aligned with the positions of NGC 253 and M82, are assigned identical weights, while the third beam direction is aligned with the position of Cen A and is weighted at 90% relative to the other two source directions. This implies that the incoming UHECRs from Cen A will be 10% fewer relative to NGC 253 and M82, respectively. Note that the monopole has been removed from the injected skymaps for the single and three-beam cases.
The choice of the injected scenarios was motivated by the observations of PAO and TA which showcases hotspots around Cen A and NGC 253 suggesting these two objects as potential UHECR sources (Abbasi et al., 2014; Aab et al., 2017; Aab et al., 2018a; Abreu et al., 2022; Abdul Halim et al., 2024). The selection of the three injected source scenarios was also motivated by theoretical models (Bell & Matthews, 2022; Taylor et al., 2023) that propose objects within the Council of Giants region as potential sources of UHECRs in our cosmic neighbourhood.
@@
here, $\ell=0....\infty$ is the multipole moment corresponding to angular scales of $180^{\circ}/\ell$, and $m$ is the azimuthal number ranging from $-\ell$ to $+\ell$. $P_{\ell{m}}(\cos\theta)$ are the associated Legendre polynomials, $\theta$ and $\phi$ are the co-latitude and azimuthal angle, and $(1/{n})(\text{d}n/\text{d}\Omega)$ represents the normalized differential cosmic ray number density per solid angle (Lang et al., 2021).
The angular power-spectrum, $C_{\ell}$, at each multipole, $\ell$, can be written in terms of the root-mean-square amplitude of the spherical harmonic coefficients by averaging over the azimuthal number ${m}$ as:
@@
Eq. 2 is used to calculate the normalised power spectra for the different injection cases considered (for the case of $L_{\rm{scat}}\approx 10$ kpc), shown in the bottom-right panel in Fig. 3. In this figure, we show the power spectrum for the extragalactic dipole scenario in green, the single beam case shown in blue, and the power spectra of the three beams scenario are shown in red. The power spectra for the injected single beam case looks smooth as the power is derived from a single spot and the rest of the map is flat. In the case of three injected beams the power spectra are more complicated with more structures on large angular scales.

### v2

@@ 1 Introduction
The arrival directions of ultra-high energy cosmic rays (UHECRs) at Earth is not entirely isotropic. In 2017, the Pierre Auger Collaboration (PAO) reported the detection of dipolar anisotropy, for UHECRs with energies $\geq 8$ EeV (Aab et al., 2017). This result was updated in Refs. Aab et al. (2018b) and Abdul Halim et al. (2024) and is now at a significance of $6.8\sigma$ with a dipole amplitude of $\delta=\rm{0.074}^{+0.010}_{-0.008}$ being obtained. The direction of this dipole lies ${\sim}~115\degree$ away from the Galactic centre, suggesting an extragalactic UHECR origin. It is worth noting that this cosmic ray dipole does not align with the CMB dipole or any potential known source.
The Telescope Array collaboration (TA) also searched for a large-scale dipole anisotropy in the northern hemisphere sky (Abbasi et al., 2020). They reported a dipole amplitude of $3.3\pm 1.9\%$ which, at their current level of statistics, is compatible in amplitude and direction with the dipole reported by PAO, but also with an isotropic distribution. When the datasets of PAO and TA are combined, a dipole with a significance of ${\sim}~4.2\sigma$ is found (Abdul Halim et al., 2023), consistent in amplitude and direction with the dipole reported by PAO alone. Besides the dipole anisotropies, several indications have been found of intermediate-scale anisotropies at higher energies as well (see e.g. Refs. Abbasi et al. (2018); Abreu et al. (2022)). However, our focus in this paper is predominantly on the large-scale anisotropy.
@@
The observed UHECR dipole strength at Earth for cosmic rays with energies above 8 EeV is at the 7% level. However, this dipole strength is not necessarily the same as the arriving extragalactic dipole amplitude, out at large distances where the influence of the Milky Way’s magnetic fields has yet to take effect. The Galactic magnetic fields can affect an external extragalactic dipole ($\delta_{\rm{EG}}$) in two main ways: the large-scale structured component changes the extragalactic direction of $\delta_{\rm{EG}}$ and the small-scale turbulent magnetic fields (with RMS value of $B_{\rm{tur}}$) suppress the strength of $\delta_{\rm{EG}}$ (Zirakashvili, 2005). Furthermore, these effects suggest that a beam (current) of particles from a source, upon propagating through the Galactic magnetic field structure, is also expected to be changed, potentially with the arrival directions being spread out by the turbulent field component, resulting in the injected power in higher multipoles migrating to lower multipoles.
The ability of the Milky Way’s turbulent magnetic field to alter the cosmic ray distribution locally observed is dependent on the spatial extent and turbulence level of this field component. Over the last few years, a growing body of evidence has accumulated indicating that a considerable level of thermal gas pressure exists out in the halo of the Milky-Way and other local galaxies of a similar mass, out to at least a distance of the virial radius (see Martynenko (2022); Bregman et al. (2022); Hopkins et al. (2022); Zhang et al. (2024)). Assuming that the plasma beta of this halo gas is close to unity, a considerable magnetic field strength is also expected out in this region (Faerman et al., 2020). Interestingly, such an inference is consistent with recent radio observations of the circumgalactic medium region (at a distance $\sim 100$ kpc) of cosmologically nearby Galaxies. From such studies, the line-of-sight magnetic field in these regions have been found to have strengths of a few hundred nG, using Faraday rotation measurements (Heesen et al., 2023).
In this work we build upon these recent results by considering the effects of a giant turbulent Galactic magnetic field halo on the penetration of cosmic rays into the Galaxy. In particular, we investigate the effect of the GMF on the arrival of an extragalactic UHECR large-scale dipole distribution, or an extragalactic UHECR small-scale localized distribution. Our focus in this paper is on the effect that turbulent GMFs have on the observed amplitude of the dipole and quadrupole.
@@ 2 Methodology - Simulation setup
UHECRs gyrating with a Larmor radius larger than the coherence length of the magnetic field, ie. $r_{L}\gg L_{\rm coh}$, in isotropic turbulent magnetic fields, undergo only small angular deflections of size $\Delta_{\Theta}=L_{\rm{coh}}/r_{L}$ in each coherent magnetic field patch, where $\Theta$ is the angle between the particle’s initial and final momentum direction. For a particle propagating a path length $L$, the number of coherence cells traversed, $N_{\rm{coh}}=L/L_{\rm{coh}}$. The standard deviation in the deflection angles from this number of small-scale deflections is,  $\sigma_{\Theta}=\sqrt{N_{\rm{coh}}}\Delta_{\Theta}$. After a sufficient number of scatterings, the spread in angles becomes large, $\sigma_{\Theta}\sim\pi$ radians. The large angle scattering length can therefore be estimated to be, $L_{\rm{scat}}\approx(\pi/\Delta_{\Theta})^{2}L_{\rm coh}=\pi^{2}\frac{r_{L}^{2}}{L_{\rm{coh}}}$. The diffusion coefficient $D\sim L_{\rm{scat}}~\rm{c}/3$, where $\rm{c}$’ is the speed of light in a vacuum. Throughout the text will use the terms ”scattering length” and ”diffusion coefficient” interchangeably.
We use CRPropa 3 (Alves Batista et al., 2016, 2022) to generate isotropic Kolmogorov turbulent magnetic fields with a power law index of -5/3 to describe the energy distribution of turbulence modes, with $L_{\rm min}$ = 450 pc and $L_{\rm max}$ = 4000 pc as the minimum and maximum wavelengths, respectively, and with a coherence length ($L_{\rm{coh}}$) of $\approx$ 1 kpc. With the coherence length of the field fixed for each case, the turbulent field effectively has only one free parameter, the magnitude of the turbulent field strength, $B_{\rm tur}$.
The turbulent fields in our setup extend out to the simulation absorption boundary (termination sphere) with radius $R_{\rm T}=60$ kpc (see Fig. 1). We note that the length scale values chosen for our setup, specifically $R_{\rm T}$ and $R_{\rm I}$, are chosen for computational rather than physical reasons. The purpose of our work, instead, is a proof of principle of the shielding effect produced by a magnetised halo, rather than an accurate description of the halo. For the giant halo scenario we consider, cosmic ray transport in the outer Galaxy is no longer 1-D, but 3D, with the cosmic rays achieving a steady state via diffusion into a larger phase space volume region, i.e. no escape boundary is needed for steady state to be achieved in this case. In this regime, the boron to carbon (B/C) ratio is no longer a probe of the true halo size ($R_{\rm T}$), but rather a probe of the size of the region within which the probability of returning back to the solar neighbourhood is of order unity (Hopkins et al., 2022, Sec. 3.1.3)).
@@ 2.1 Estimation of the scattering length
[…]
Figure 1: Schematic representation of the simulation setup for the dipole and beam setup shown on the left and right-hand side, respectively. In both figures, $R_{\rm T}$, $R_{\rm O}$ and $R_{\rm I}$ are the radii of the termination, observer and injection spheres, respectively. The changing colour of the green ring from right to left indicates the changing strength of the external dipole magnitude, and the single arrow in the beam setup denotes injected UHECRs with the arrow directions pointing inwards in both setups. For the extragalactic dipole case, we adopted a Lambert distribution of particle directions.
Comparing these numerical results with our earlier analytic expectations, for small values of $B_{\rm{tur}}$, such that $r_{L}\gg L_{\rm coh}$, the relation $L_{\rm{scat}}\approx\pi^{2}\frac{r_{L}^{2}}{L_{\rm{coh}}}$ appears to approximately hold (see Fig. 2). This dependence of the scattering length on the particle Larmor radius, however, is expected to change once the magnetic field strength is further increased, and the UHECR Larmor radius drops below the magnetic field coherence length.
For the case when the Larmor radius is below the coherence length scale, $r_{L}<L_{\rm{coh}}$, resonant interactions with the turbulent modes become the dominant interaction of the UHECR with the turbulence. The transition to this regime is indicated by the vertical black line in Fig. 2. Due to the limited turbulence dynamical range that we simulate, this regime cannot be strongly probed by our simulations.
Figure 2: The scattering length ($\star$) plotted as a function of the Larmor radius on the bottom x-axis, and the corresponding turbulent magnetic field strength on the top-secondary x-axis, for 8.5 EeV protons. The green dashed line shows the expected results applicable when $r_{L}>L_{\rm coh}$. Left of the dashed vertical black line shows the region where $r_{\rm{L}}<L_{\rm{coh}}$ and deviation from the small angle scattering regime is expected.
@@ 2.2 Calculation of UHECR dipole amplitudes through turbulent magnetic fields
In order to create synthetic skymaps for an observer we inject $N_{\rm{inj}}\approx 10^{10}$ UHECRs into the setup depicted in Fig. 1. This setup ensures that the observed skymaps contain good statistics ($N_{\rm{obs}}\approx 4\times 10^{5}$). In reality, UHECRs enter the Galaxy from all directions, and what is observed by PAO (Aab et al., 2017; Abreu et al., 2022) is a fraction of the UHECRs that hit the observer (Earth).
Figure 3: Top-left: panel shows the injected extragalactic dipole skymap. Top-right: shows the single-beam injected skymap. Bottom-left: panel depicts the combined three-beam injected skymap. Bottom-right: panel shows the comparative power spectrum analysis of the three injected skymaps. Note that the maps shown here have a $10^{\circ}$ top-hat smoothing applied.
Forward tracking of UHECRs was performed using CRPropa 3, where 8.5 EeV protons were propagated through the Galaxy’s turbulent magnetic field component after being injected inward from an injection sphere. The observer sphere radius $R_{\rm O}$ was set to 1 kpc, comparable to the size of $L_{\rm{coh}}$. The simulation ends when the particle reaches either the observer sphere or the termination sphere (see Fig. 1). The injection sphere was kept at a radius, $R_{\rm I}=$ 30 kpc, and the escape boundary was twice as large as the injection sphere at $R_{\rm T}=$ 60 kpc.
@@
    An extragalactic dipole case where we impose an external dipole at $\sqrt{3}$ level at the injection sphere (see top-left skymap Fig. 3). The value $\sqrt{3}$ comes from the prefactor in $\delta_{0}=\sqrt{3}\frac{C_{1}}{C_{0}}$. We adopted a Lambert distribution (isotropic) for every injection point on the injection sphere.
@@
    A three beam case where we inject three $10^{\circ}$ beams of UHECRs at the positions of Cen A, NGC 253 ($l=97.4^{\circ}$, $b=-87.9^{\circ}$) and M82 ($l=141.4^{\circ}$, $b=+40.6^{\circ}$) (see bottom-left skymap Fig. 3). In the three-beam configuration, two of the beam directions, aligned with the positions of NGC 253 and M82, are assigned identical weights of 35%, while the third beam direction, aligned with the position of Cen A, is weighted at 30%. Note that the monopole has been removed from the injected skymaps for the single and three-beam cases. Additionally, we smooth the injected skymaps by $10^{\circ}$.
The choice of the injected scenarios was motivated by the observations of PAO and TA which showcases hotspots around Cen A and NGC 253 suggesting these two objects as potential UHECR sources (Abbasi et al., 2014; Aab et al., 2017, 2018a; Abreu et al., 2022; Abdul Halim et al., 2024). The selection of the three injected source scenarios was also motivated by theoretical models (Bell & Matthews, 2022; Taylor et al., 2023) that propose objects within the Council of Giants region as potential sources of UHECRs in our cosmic neighbourhood.
@@
here, $\ell=0.$$\infty$ is the multipole moment corresponding to angular scales of $180^{\circ}/\ell$, and $m$ is the azimuthal number ranging from -$\ell$ to $\ell$. $P_{\ell{m}}(\cos\theta)$ are the associated Legendre polynomials, $\theta$ and $\phi$ are the co-latitude and azimuthal angle, and $(1/{n})(\text{d}n/\text{d}\Omega)$ represents the normalized differential cosmic ray number density per solid angle (Lang et al., 2021).
The angular power-spectrum, $C_{\ell}$, at each multipole, $\ell$, can be written in terms of the root-mean-square amplitude of the spherical harmonic coefficients by averaging over the azimuthal number m{m} as:
@@
Eq. 2 is used to calculate the normalised power spectra for the different injection cases considered (for the case of $L_{\rm{scat}}\approx 80$ kpc), shown in the bottom-right panel in Fig. 3. In this figure, we show the power spectrum for the extragalactic dipole scenario in green, the single beam case shown in blue, and the power spectrum of the three beams scenario is shown in red. The power spectra for the injected dipole case shows that the power resides in the monopole and the dipole components. The higher order multipoles are at the noise level. The power spectra for the injected single beam case looks smooth as the power is derived from a single spot and the rest of the map is flat. In the case of three injected beams the power spectra are more complicated with more structures on large angular scales.